# YOLO Evolution: A Comprehensive Benchmark and Architectural Review of YOLOv12, YOLO11, and Their Previous Versions

### v3

@@ III-A2 Africa Wild Life Dataset
III-A3 Ships/Vessels Dataset
The Ships/Vessels dataset is an extensive open-source collection containing approximately 13.5k images, collected by Siddharth Sah from numerous Roboflow datasets, curated explicitly for ship detection [55]. Each image has been manually annotated with bounding boxes in the YOLO format, ensuring precise and efficient detection of ships. This dataset features a single class, ”ship,” and is divided into 70% training, 20% validation, and 10% testing allowing for streamlined and focused analysis. Notably, no data augmentation techniques were applied during the experiment, which ensures that the model’s performance is evaluated on the raw, unaltered dataset. However, the relatively small size of the target objects and their varying rotations pose challenges for detection, particularly for the YOLO algorithm, which often struggles with small object detection and objects with varying orientations. The dataset is essential for various practical applications such as maritime safety, fisheries management, marine pollution monitoring, defense, maritime security, and more.
III-B Models
III-B1 Ultralytics vs. Original YOLO
In this subsection, we will conduct a comparative analysis between the models provided by Ultralytics and their original counterparts on the Traffic Signs dataset provided by Radu Oprea [43] using the same hyperparameters in Table V. The objective is to highlight the differences between Ultralytics models and the original versions, which justifies the exclusion of YOLOv4 [7], YOLOv6 [31], and YOLOv7 [69] from this paper due to the lack of support for these models by Ultralytics. This analysis will demonstrate why focusing exclusively on Ultralytics-supported models ensures a fair and consistent benchmark evaluation.
@@ III-B Models
Ultralytics library provides researchers and programmers with various YOLO models for inference, validation, training, and export. Based on the results of Table I, we notice that Ultralytics does not support YOLOv1, YOLOv2, YOLOv4, and YOLOv7. Concerning YOLOv6, the library only supports the configuration *.yaml files without the pre-trained *.pt models.
TABLE I: Ultralytics-supported library tasks and models
@@
Based on the results of our comparative analysis of the Ultralytics models and their original counterparts on the Traffic Signs dataset presented in Table II, we observe significant discrepancies between the performance of the Ultralytics models and their original counterparts. Notably, Ultralytics’ versions of YOLOv5n (nano) and YOLOv3 demonstrate superior performance, underscoring the enhancements and optimizations implemented by Ultralytics. Conversely, the original YOLOv9c (compact) slightly outperforms its Ultralytics version, potentially due to the lack of extensive optimization for this newer model by Ultralytics. These observations highlight that the Ultralytics models have undergone substantial modifications, making a direct comparison with the original versions inequitable. Consequently, the noticeable performance discrepancy between the two models, including the original and Ultralytics models in the same benchmarking study, would not provide a fair or accurate assessment. Therefore, this paper will focus exclusively on the Ultralytics-supported versions to ensure consistent and fair benchmarks.
In total, 33 models from 7 different YOLO versions were trained on 3 different datasets as seen in table III. Each YOLO version includes several scaled models, (e.g., YOLOv5u, YOLOv5un, YOLOv5us, and YOLOv5ux), with the suffixes denoting model size and complexity, such as ”n” for nano, ”s” for small, ”m” for medium, ”l” for large, ”x” for extra-large, ”t” for tiny, ”c” for compact, ”b” for balanced, and ”e” for extended. These models offer a variety of trade-offs between accuracy and inference speed, as discussed in the following sections.
TABLE II: Ultralytics and original YOLO performance comparison
@@
TABLE III: YOLO versions and scaled versions
Version Scaled Version
YOLOv3u YOLOv3u-tiny
YOLOv3u
YOLOv5u YOLOv5un (nano)
YOLOv5us (small)
YOLOv5um (medium)
YOLOv5ul (large)
YOLOv5ux (extra-large)
YOLOv8 YOLOv8n (nano)
YOLOv8s (small)
YOLOv8m (medium)
YOLOv8l (large)
YOLOv8x (extra-large)
YOLOv9 YOLOv9t (tiny)
YOLOv9s (small)
YOLOv9m (medium)
YOLOv9c (compact)
YOLOv9e (extended)
YOLOv10 YOLOv10n (nano)
YOLOv10s (small)
YOLOv10m (medium)
YOLOv10b (balanced)
YOLOv10l (large)
YOLOv10x (extra-large)
YOLO11 YOLO11n (nano)
YOLO11s (small)
YOLO11m (medium)
YOLO11l (large)
YOLO11x (extra-large)
YOLOv12 YOLO11n (nano)
YOLOv12s (small)
YOLOv12m (medium)
YOLOv12l (large)
YOLOv12x (extra-large)
III-B2 YOLOv3u
@@
YOLOv3, based on its predecessors, aims to improve localization errors and detection efficiency, particularly for smaller objects. It uses the Darknet-53 framework, which has 53 convolutional layers and achieves double the speed of ResNet-152 [52]. YOLOv3 also incorporates elements from the Feature Pyramid Network (FPN), such as residual blocks, skip connections, and up-sampling, to enhance its ability to detect objects efficiently across varying scales, as seen in Figure 4. The algorithm generates feature maps at three distinct scales, down-sampling the input at factors of 32, 16, and 8, and uses a three-scale detection mechanism to detect large, medium, and small-sized objects using distinct feature maps. Despite its improvements, YOLOv3 still faces challenges in achieving precise results for medium and large-sized objects, so Ultralytics released YOLOv3u. YOLOv3u is an improved version of YOLOv3 that utilizes an anchor-free detection method used later in YOLOv8 and improves upon the accuracy and speed of YOLOv3, especially for medium and large-sized objects.
@@
Figure 4: YOLOv3 architecture showcasing the residual blocks and the upsampling layers [10].
III-B3 YOLOv5u
YOLOv5, proposed by Glenn Jocher, transitions from the Darknet framework to PyTorch, retaining many improvements from YOLOv4 [64] [26] and utilizing CSPDarknet as its backbone. CSPDarknet is a modified version of the original Darknet architecture that incorporates Cross-Stage Partial connections by splitting feature maps into separate paths, allowing for more efficient feature extraction and reduced computational costs. YOLOv5 features a strided convolution layer with a large window size, aiming to reduce memory and computational costs, as showcased in Figure 5. Moreover, this version adopts the Spatial Pyramid Pooling Fast (SPPF) module to provide a multiscale representation of the input feature maps. The SPPF module works by pooling features at different scales and concatenating them, allowing the network to capture fine and coarse information. This helps recognize objects of various sizes more effectively. In addition, YOLOv5 implements several augmentations, such as Mosaic, copy-paste, random affine, MixUp, HSV augmentation, and random horizontal flip. YOLOv5 is available in five variants, varying in width and depth of convolution modules. Ultralytics is actively improving this model through YOLOv5u, adopting an anchor-free detection method and achieving better overall performance, especially on complex objects of different sizes.
@@ III-B3 YOLOv5u
Figure 5: Detailed architecture of YOLOv5 including the CSPDarknet Backbone, PANet Neck, and YOLO Layer Head [14].
III-B4 YOLOv8
@@
Ultralytics has introduced YOLOv8, a significant evolution in the YOLO series, with five scaled versions [61] [27]. Alongside object detection, YOLOv8 also provides various applications such as image classification, pose estimation, instance segmentation, and oriented object detection (OOB). Key features include a backbone similar to YOLOv5, with adjustments in the CSPLayer, now known as the C2f module, which combines high-level features with contextual information for enhanced detection accuracy highlighted in Figure 6. YOLOv8 also introduces a semantic segmentation model called YOLOv8-Seg, which combines a CSPDarknet53 feature extractor with a C2F module, achieving state-of-the-art results in object detection and semantic segmentation benchmarks while maintaining high efficiency.
@@
Figure 6: Detailed architecture of YOLOv8 showcasing the backbone’s multiple convolutional layers to extract hierarchical features, the Feature Pyramid Network (FPN) enhances detection at different scales. The network head performs final predictions, incorporating convolutional blocks and upsample blocks to refine features [30].
III-B5 YOLOv9
YOLOv9, developed by Chien-Yao Wang, I-Hau Yeh, and Hong-Yuan Mark Liao, uses the Information Bottleneck Principle and Reversible Functions to preserve essential data across the network’s depth, ensuring reliable gradient generation and improved model convergence and performance [70]. Reversible functions, which can be inverted without loss of information, are another cornerstone of YOLOv9’s architecture. This property allows the network to retain a complete information flow, enabling more accurate updates to the model’s parameters. Moreover, YOLOv9 offers five scaled versions for different uses, focusing on lightweight models, which are often under-parameterized and prone to losing significant information during the feedforward process.
Programmable Gradient Information (PGI) is a significant advancement introduced in YOLOv9. PGI is a method that dynamically adjusts the gradient information during training to optimize learning efficiency. By selectively focusing on the most informative gradients, PGI helps preserve crucial information that might be lost in lightweight models. This advancement ensures the model retains the essential features for accurate object detection, improving overall performance.
In addition, YOLOv9 incorporates Gradient Enhanced Lightweight Architecture Network (GELAN), a new architectural advancement designed to improve parameter utilization and computational efficiency as illustrated in Figure 7. GELAN achieves this by optimizing the computational pathways within the network, allowing for better resource management and adaptability to various applications without compromising speed or accuracy.
@@ III-B5 YOLOv9
Figure 7: YOLOv9 architecture featuring CSPNet, ELAN, and GELAN modules. CSPNet optimizes gradient flow and reduces computational complexity via feature map partitioning. ELAN enhances learning efficiency by linearly aggregating features, and GELAN extends this concept by integrating features from various depths and pathways [70].
III-B6 YOLOv10
@@
YOLOv10, developed by Tsinghua University researchers, builds upon previous models’ strengths with key innovations [68]. The architecture has an enhanced Cross Stage Partial Network (CSPNet) backbone for improved gradient flow and reduced computational redundancy, as seen in Figure 8. The network is structured into three main components: the backbone, the neck, and the detection head. The neck includes Path Aggregation Network (PAN) layers for effective multiscale feature fusion. PAN is designed to enhance information flow by aggregating features from different layers, enabling the network to better capture and combine details at various scales, which is crucial for detecting objects of different sizes. At the same time, the One-to-Many Head generates multiple predictions per object during training to provide rich supervisory signals and improve learning accuracy. Five scaled versions are available, from nano to extra-large.
@@
For inference, the One-to-One Head generates a single best prediction per object, eliminating the need for Non-Maximum Suppression (NMS). By removing the need for NMS, YOLOv10 reduces latency and improves the post-processing speed. In addition, YOLOv10 includes NMS-Free Training, which uses consistent dual assignments to reduce inference latency, and a model design that optimizes various components from both efficiency and accuracy perspectives. This includes lightweight classification heads, spatial-channel decoupled downsampling, and rank-guided block design. In addition, the model incorporates large-kernel convolutions and partial self-attention modules to enhance performance without significant computational costs.
@@
Figure 8: YOLOv10 architecture showcasing the dual label assignment strategy for improving accuracy. The PAN layer enhances feature representation alongside one-to-many head for regression and classification tasks and one-to-one head for precise localization [68].
III-B7 YOLO11
@@
YOLO11 [28] is one of the latest additions to the YOLO series developed by Ultralytics, building upon the developments of its predecessors, especially YOLOv8. This iteration offers five scaled models from nano to extra large, catering to various applications. Like YOLOv8, YOLO11 includes numerous applications such as object detection, instance segmentation, image classification, pose estimation, and oriented object detection (OBB).
@@
Figure 9: YOLO11 architecture showcasing the new C3k2 blocks and the C2PSA module. [28] [46].
III-B8 YOLOv12
YOLOv12 [63] is the latest evolution in the YOLO series, introducing an attention-centric design that significantly enhances both speed and accuracy. It continues the trend of offering five scalable models (Nano to Extra Large), making it adaptable for a wide range of applications such as object detection, instance segmentation, and oriented object detection (OBB). As seen in Figure 10, YOLOv12 introduces the Area Attention (A2) module, which maintains a large receptive field while drastically reducing computational complexity, allowing the model to enhance speed without compromising accuracy. Additionally, it features Residual Efficient Layer Aggregation Networks (R-ELAN), which improve training stability and model convergence through block-level residual design and optimized feature aggregation.
@@ III-B8 YOLOv12
Figure 10: YOLOv12 architecture showcasing the new Area Attention (A2) module and Residual Efficient Layer Aggregation Networks (R-ELAN). [63]
III-C Hardware and Software Setup
Table IV presents the libraries and software versions used throughout this study. The experiments were conducted with Python 3.12, Ubuntu 22.04, CUDA 12.5, and cuDNN 8.9.7 for GPU acceleration. Ultralytics 8.2.55 was used for model training, while WandB 0.17.4 was utilized for tracking experiments. To ensure a fair comparison, similar hyperparameters were used across all models, as outlined in Table V. The key hyperparameters include 100 epochs, the AdamW optimizer, a batch size of 16, and an image size of (640, 640). The learning rate was initially set to 0.0001 and increased to 0.01, while the dropout rate was set to 0.15. The dataset was split into 70% for training, 20% for validation, and 10% for testing with no data augmentation techniques applied to any dataset. The experiments were carried out using 2 NVIDIA RTX 4090 GPUs, each with 16,384 CUDA cores.
TABLE IV: Software setup
Name Version
Python 3.12
Ubuntu 22.04
CUDA 12.5
cuDNN 8.9.7
Ultralytics 8.2.55
WandB 0.17.4
TABLE V: Table of parameters
@@ III-D Metrics
This study evaluates the performance of YOLO models using three primary metrics: accuracy, computational efficiency, and size. The accuracy metrics include Precision, Recall, Mean Average Precision at an Intersection over Union (IoU) threshold of 0.50 (mAP50), and Mean Average Precision at an Intersection over Union (IoU) threshold of 0.50-0.95 (mAP50-95). Precision [44] measures the ratio of correctly predicted observations to the total predicted observations, thus highlighting the occurrence of False Positives. Conversely, Recall [44] measures the ratio of correctly predicted observations to all actual observations, thus emphasizing the occurrence of False Negatives. Both mAP50 and mAP50-95 [75] provide a comprehensive summary of Precision and Recall. While mAP50 calculates the Mean Average Precision at an IoU threshold of 0.50, mAP50-95 extends this calculation across multiple IoU thresholds from 0.50 to 0.95, with 0.05 step size.
Regarding computational efficiency metrics, Preprocessing Time, Inference Time, and Postprocessing Time will be utilized to evaluate the model’s speed. Preprocessing Time refers to the duration taken to prepare raw data for input into the model. Inference Time is the duration required for the model to process the input data and generate predictions. Postprocessing Time denotes the time needed to convert the model’s raw predictions into a final, usable format. These metrics were measured using a sample of images for testing after training the models. Additionally, the Giga Floating-Point Operations Per Second (GFLOPs) measure the computational power for the model training, reflecting its efficiency. In contrast, the size metric reflects the actual disk size of the model and the number of its parameters.
These metrics are essential for providing a comprehensive overview of YOLO models’ performance, allowing for effective comparison and evaluation. These metrics allows a thorough assessment of the accuracy and efficiency of different YOLO models, ensuring a robust benchmark for their performance and applications in various real-world scenarios.
IV Benchmark Results and Discussion
IV-A Results
IV-A1 Traffic Signs Dataset
TABLE VI: Evaluation results for the Traffic Signs Dataset.
Versions Precision Recall mAP50 mAP50-95
Preprocess
Time
Inference
Time
Postprocess
Time
Total
Time
GFLOPs Size
YOLOv3u 0.75 0.849 0.874 0.781 0.7 8.5 0.4 9.6 207.86 282.4
YOLOV3u tiny 0.845 0.667 0.772 0.682 1.4 0.7 0.3 2.4 24.44 19
YOLOv5un 0.805 0.679 0.749 0.665 0.6 6.6 0.4 7.6 5.65 7.1
YOLOv5us 0.85 0.777 0.827 0.744 0.5 7.8 0.4 8.7 18.58 23.9
YOLOv5um 0.849 0.701 0.83 0.744 1.1 9.5 0.4 11 50.54 64.1
YOLOv5ul 0.831 0.836 0.886 0.799 0.6 9.7 0.4 10.7 106.85 134.9
YOLOv5ux 0.863 0.795 0.867 0.777 1.1 9.8 0.4 11.3 195.2 246.3
YOLOv8n 0.749 0.688 0.777 0.689 0.6 6.8 0.4 7.8 6.55 8.1
YOLOv8s 0.766 0.788 0.806 0.718 0.6 7.8 0.4 8.8 22.59 28.6
YOLOv8m 0.838 0.805 0.845 0.763 1.6 9.1 0.4 11.1 52.12 78.9
YOLOv8l 0.771 0.789 0.853 0.767 0.6 9.2 0.4 10.2 87.77 165
YOLOv8x 0.902 0.744 0.874 0.78 0.6 9.4 0.4 10.4 136.9 257.7
YOLOv9t 0.792 0.748 0.812 0.731 0.5 10 0.4 10.9 4.93 7.7
YOLOv9s 0.763 0.81 0.828 0.75 0.6 11.1 0.4 12.1 15.33 26.8
YOLOv9m 0.864 0.796 0.864 0.784 1 12.1 0.4 13.5 40.98 76.7
YOLOv9c 0.827 0.807 0.852 0.769 1.3 11.6 0.4 13.3 51.8 102.6
YOLOv9e 0.819 0.824 0.854 0.764 0.8 16.1 0.4 17.3 117.5 189.4
YOLOv10n 0.722 0.602 0.722 0.64 1 0.8 0.2 2 5.59 8.3
YOLOv10s 0.823 0.742 0.834 0.744 1.2 1.1 0.2 2.5 15.9 24.7
YOLOv10m 0.834 0.843 0.88 0.781 1.2 2.4 0.2 3.8 32.1 63.8
YOLOv10b 0.836 0.764 0.859 0.765 1 3.1 0.2 4.3 39.7 98.4
YOLOv10l 0.873 0.807 0.866 0.771 1.1 3.8 0.2 5.1 50 126.8
YOLOv10x 0.773 0.854 0.88 0.787 1 6.3 0.2 7.5 61.4 170.4
YOLO11n 0.768 0.695 0.757 0.668 1.2 0.6 0.4 2.2 5.35 6.4
YOLO11s 0.819 0.758 0.838 0.742 1.2 1 0.4 2.6 18.4 21.4
YOLO11m 0.898 0.826 0.893 0.795 1.2 2.4 0.4 4 38.8 67.9
YOLO11l 0.862 0.839 0.889 0.794 1.2 3 0.4 4.6 49 86.8
YOLO11x 0.819 0.816 0.885 0.784 0.9 6.1 0.4 7.4 109 194.8
YOLOv12n 0.883 0.684 0.802 0.708 1.2 4.6 0.4 6.2 5.34 6.4
YOLOv12s 0.761 0.839 0.877 0.78 1.6 4.3 0.5 6.4 18.1 21.4
YOLOv12m 0.868 0.811 0.89 0.79 0.9 4.9 0.4 6.2 39 67.3
YOLOv12l 0.826 0.836 0.877 0.781 0.9 5.5 0.4 6.8 51.2 89.6
YOLOv12x 0.832 0.849 0.889 0.793 1 9.5 0.5 11 114 198.9

### v4

@@ 3.1.2 Africa Wild Life Dataset
3.1.3 Ships/Vessels Dataset
The Ships/Vessels dataset is an extensive open-source collection containing approximately 13.5k images, collected by Siddharth Sah from numerous Roboflow datasets, curated explicitly for ship detection [65]. Each image has been manually annotated with bounding boxes in the YOLO format, ensuring precise and efficient detection of ships. This dataset features a single class, ”ship,” and is divided into 70% training, 20% validation, and 10% testing. Notably, no data augmentation techniques were applied during the experiment, which ensures that the model’s performance is evaluated on the raw, unaltered dataset. However, the relatively small size of the target objects, varying aspect ratios, and their varying rotations pose challenges for detection, particularly for the YOLO algorithm, which often struggles with small object detection and objects with varying orientations. The dataset is essential for various practical applications such as maritime safety, fisheries management, marine pollution monitoring, defense, maritime security, and more.
3.2 Models
3.2.1 Ultralytics vs. Original YOLO
In this subsection, we will conduct a comparative analysis between the models provided by Ultralytics and their original counterparts on the Traffic Signs dataset provided by Radu Oprea [51] using the same hyperparameters in Table 3. The objective is to highlight the differences between Ultralytics models and the original versions, which justifies the exclusion of YOLOv4 [10], YOLOv6 [38], and YOLOv7 [81] from this paper due to the lack of support for these models by Ultralytics. This analysis will demonstrate why focusing exclusively on Ultralytics-supported models ensures a fair and consistent benchmark evaluation.
@@ 3.2 Models
Ultralytics library provides researchers and programmers various YOLO models for inference, validation, training, and export. Based on the results of Table 1, we notice that Ultralytics does not support YOLOv1, YOLOv2, YOLOv4, and YOLOv7. Concerning YOLOv6, the library only supports the configuration *.yaml files without the pre-trained *.pt models.
Table 1: Ultralytics-supported library tasks and models
@@
Based on the results of our comparative analysis of the Ultralytics models and their original counterparts on the Traffic Signs dataset presented in Table 2, we observe significant discrepancies between the performance of the Ultralytics models and their original counterparts. Notably, Ultralytics’ versions of YOLOv5n (nano) and YOLOv3 demonstrate superior performance, underscoring the enhancements and optimizations implemented by Ultralytics. Conversely, the original YOLOv9c (compact) slightly outperforms its Ultralytics version, potentially due to the lack of extensive optimization for this newer model by Ultralytics. These observations highlight that the Ultralytics models have undergone substantial modifications, making a direct comparison with the original versions inequitable. Consequently, the noticeable performance discrepancy between the two models, including the original and Ultralytics models in the same benchmarking study, would not provide a fair or accurate assessment. Therefore, this paper will focus exclusively on the Ultralytics-supported versions to ensure consistent and fair benchmarks.
In total, 33 models from 7 different YOLO versions were trained on three different datasets, as seen in Figure 4. Each YOLO version includes several scaled models (e.g., YOLOv5u, YOLOv5un, YOLOv5us, and YOLOv5ux), with the suffixes denoting model size and complexity, such as ”n” for nano, ”s” for small, ”m” for medium, ”l” for large, ”x” for extra-large, ”t” for tiny, ”c” for compact, ”b” for balanced, and ”e” for extended. These models offer a variety of trade-offs between accuracy and inference speed, as discussed in the following sections.
Table 2: Ultralytics and original YOLO performance comparison
@@
3.2.2 YOLOv3u
@@
YOLOv3 enhances localization and detection efficiency, particularly for small objects, using the Darknet-53 framework, which offers double the speed of ResNet-152 [61]. It integrates Feature Pyramid Network (FPN) elements, including residual blocks, skip connections, and up-sampling, to improve multi-scale detection, as illustrated in Figure 5. The model generates feature maps at three scales (down-sampling at factors of 32, 16, and 8) for detecting objects of varying sizes. However, YOLOv3 struggles with medium and large objects, leading Ultralytics to introduce YOLOv3u, which adopts an anchor-free detection method, later used in YOLOv8, improving both accuracy and speed.
@@
Figure 5: YOLOv3 architecture showcasing the residual blocks and the upsampling layers [13]
3.2.3 YOLOv5u
@@
Figure 6: Detailed architecture of YOLOv5 including the CSPDarknet Backbone, PANet Neck, and YOLO Layer Head [17]
YOLOv5, developed by Glenn Jocher, transitions from Darknet to PyTorch, incorporating CSPDarknet as its backbone for improved feature extraction and reduced computational cost [74] [31]. It features Cross-Stage Partial (CSP) connections, a strided convolution layer, and the Spatial Pyramid Pooling Fast (SPPF) module, which enhances multi-scale feature representation, as illustrated in Figure 6. YOLOv5 also integrates data augmentation techniques (Mosaic, copy-paste, MixUp, HSV augmentation) to improve robustness. Available in five scaled variants, it continues to evolve with YOLOv5u, which introduces anchor-free detection for enhanced accuracy and efficiency, particularly on complex objects of varying sizes.
3.2.4 YOLOv8
@@
Ultralytics has introduced YOLOv8, a significant evolution in the YOLO series, with five scaled versions [71] [32]. Alongside object detection, YOLOv8 also provides various applications such as image classification, pose estimation, instance segmentation, and OBB. Key features include a backbone similar to YOLOv5, with adjustments in the CSPLayer, now known as the C2f module, which combines high-level features with contextual information for enhanced detection accuracy highlighted in Figure 7. YOLOv8 also introduces a semantic segmentation model called YOLOv8-Seg, which combines a CSPDarknet53 feature extractor with a C2F module, achieving state-of-the-art results in object detection and semantic segmentation benchmarks while maintaining high efficiency.
@@
Figure 7: Detailed architecture of YOLOv8 showcasing the backbone’s multiple convolutional layers to extract hierarchical features, the Feature Pyramid Network (FPN) enhances detection at different scales. The network head performs final predictions, incorporating convolutional blocks and upsample blocks to refine features [35]
3.2.5 YOLOv9
@@
Figure 8: YOLOv9 architecture featuring CSPNet, ELAN, and GELAN modules. CSPNet optimizes gradient flow and reduces computational complexity via feature map partitioning. ELAN enhances learning efficiency by linearly aggregating features, and GELAN extends this concept by integrating features from various depths and pathways [82]
YOLOv9, developed by Chien-Yao Wang, I-Hau Yeh, and Hong-Yuan Mark Liao, enhances object detection using the Information Bottleneck Principle and Reversible Functions, ensuring efficient gradient propagation and improved model convergence [82]. Reversible functions enable lossless information retention, benefiting lightweight models prone to under-parameterization. A key advancement, Programmable Gradient Information (PGI), dynamically adjusts gradient flow during training, prioritizing informative gradients to prevent feature loss. Additionally, YOLOv9 integrates Gradient Enhanced Lightweight Architecture Network (GELAN), optimizing computational pathways for better parameter utilization, high-speed inference, and accurate detection, as illustrated in Figure 8. With five scalable versions, YOLOv9 balances efficiency and feature retention, making it highly adaptable for various real-world applications.
3.2.6 YOLOv10
@@
YOLOv10, developed by Tsinghua University researchers, introduces key innovations, enhancing gradient flow and computational efficiency through an improved Cross Stage Partial Network (CSPNet) backbone [79]. The architecture, illustrated in Figure 9, consists of three main components: the backbone, the neck, and the detection head. The Path Aggregation Network (PAN) neck improves multi-scale feature fusion, allowing for better detection of objects at different sizes. The One-to-Many Head generates multiple predictions per object during training, enhancing learning accuracy.
@@
For inference, YOLOv10 replaces Non-Maximum Suppression (NMS) with a One-to-One Head, producing a single best prediction per object, reducing latency and post-processing time. Additionally, NMS-Free Training leverages dual assignments, optimizing both speed and accuracy. Other enhancements include lightweight classification heads, spatial-channel decoupled downsampling, large-kernel convolutions, and partial self-attention modules, all improving efficiency without significantly increasing computational costs. YOLOv10 is available in five scaled versions, from nano to extra-large, catering to diverse use cases.
@@
Figure 9: YOLOv10 architecture showcasing the dual label assignment strategy for improving accuracy. The PAN layer enhances feature representation alongside one-to-many head for regression and classification tasks and one-to-one head for precise localization [79]
3.2.7 YOLO11
Figure 10: YOLO11 architecture showcasing the new C3k2 blocks and the C2PSA module [33] [54]
@@ 3.2.7 YOLO11
YOLO11 [33] is one of the latest additions to the YOLO series developed by Ultralytics, building upon the developments of its predecessors, especially YOLOv8. This iteration offers five scaled models from nano to extra large, catering to various applications. Like YOLOv8, YOLO11 includes numerous applications such as object detection, instance segmentation, image classification, pose estimation, and OBB.
@@
3.2.8 YOLOv12
Figure 11: YOLOv12 architecture showcasing the new Area Attention (A2) module and Residual Efficient Layer Aggregation Networks (R-ELAN)
YOLOv12 [73] is the latest evolution in the YOLO series, introducing an attention-centric design that significantly enhances both speed and accuracy. It continues the trend of offering five scalable models (Nano to Extra Large), making it adaptable for a wide range of applications such as object detection, instance segmentation, and OBB. As seen in Figure 11, YOLOv12 introduces the Area Attention (A2) module, which maintains a large receptive field while drastically reducing computational complexity, allowing the model to enhance speed without compromising accuracy. Additionally, it features Residual Efficient Layer Aggregation Networks (R-ELAN), which improve training stability and model convergence through block-level residual design and optimized feature aggregation.
@@ 3.2.8 YOLOv12
3.3 Hardware and Software Setup
The experiments were conducted with Python 3.12, Ubuntu 22.04, CUDA 12.4, and cuDNN 8.9.7 for GPU acceleration. Ultralytics 8.2.55 was used for model training, while WandB 0.17.4 was utilized for tracking experiments. To ensure a fair comparison, similar hyperparameters were used across all models, as outlined in Table 3. The experiments were carried out using 2 NVIDIA RTX 4090 GPUs, each with 16,384 CUDA cores.
Table 3: Table of parameters
@@ 3.4 Metrics
This study evaluates YOLO models based on accuracy, computational efficiency, and size. Accuracy metrics include Precision, Recall, mAP50, and mAP50-95. Precision measures the ratio of correctly predicted observations to total predictions, indicating False Positives, while Recall highlights False Negatives by comparing correct predictions to actual observations [52]. mAP50 calculates Mean Average Precision at an IoU threshold of 0.50, whereas mAP50-95 extends this across thresholds from 0.50 to 0.95 with a 0.05 step size [88].
Computational efficiency is assessed using Preprocessing Time, Inference Time, and Postprocessing Time. Preprocessing Time refers to data preparation, Inference Time measures prediction generation, and Postprocessing Time converts raw outputs into final results. Additionally, GFLOPs indicate computational power, while model size reflects storage requirements and parameter count. These metrics enable a comprehensive performance comparison, ensuring a robust benchmark for real-world applications.
4 Benchmark Results
4.1 Results
4.1.1 Traffic Signs Dataset
Figure 12: Evaluation results for the traffic signs dataset.